Micah needs $18,500 to purchase a boat. How much money will he need to invest today in a savings account earning 3.2% interest, compounding quarterly, to have enough money to purchase the boat in 14 years?
Answer by Alan3354(69443)

Micah needs $18,500 to purchase a boat. How much money will he need to invest today in a savings account earning 3.2% interest, compounding quarterly, to have enough money to purchase the boat in 14 years?
---------------
Micah's really planning ahead.
FV = PV*(1 + r)^t
3.2% annually --> 0.8% per quarter
14 years = 56 quarters
18500 = PV*(1 + 0.008)^56
PV = 18500/(1.008^56)
PV =~ $11840.82